# Learning Temporal 3D Semantic Scene Completion via Optical Flow Guidance

### v1

@@ 3.5 Occlusion-Guided Voxel Refinement
Methods Venues Input IoU
[…]
mIoU
MonoScene [3] CVPR’2022 S 34.16 54.70 27.10 24.80 5.70 14.40 18.80 3.30 0.50 0.70 4.40 14.90 2.40 19.50 1.00 1.40 0.40 11.10 3.30 2.10 11.08
TPVFormer [10] CVPR’2023 S 34.25 55.10 27.20 27.40 6.50 14.80 19.20 3.70 1.00 0.50 2.30 13.90 2.60 20.40 1.10 2.40 0.30 11.00 2.90 1.50 11.26
OccFormer [50] ICCV’2023 S 34.53 55.90 30.30 31.50 6.50 15.70 21.60 1.20 1.50 1.70 3.20 16.80 3.90 21.30 2.20 1.10 0.20 11.90 3.80 3.70 12.32
Symphonize [13] CVPR’2024 S 42.19 58.40 29.30 26.90 11.70 24.70 23.60 3.20 3.60 2.60 5.60 24.20 10.00 23.10 3.20 1.90 2.00 16.10 7.70 8.00 15.04
BRGScene [14] IJCAI’2024 S 43.34 61.90 31.20 30.70 10.70 24.20 22.80 2.80 3.40 2.40 6.10 23.80 8.40 27.00 2.90 2.20 0.50 16.50 7.00 7.20 15.36
CGFormer [47] NIPS’2024 S 44.41 64.30 34.20 34.10 12.10 25.80 26.10 4.30 3.70 1.30 2.70 24.50 11.20 29.30 1.70 3.60 0.40 18.70 8.70 9.30 16.63
VoxFormer-T [20] CVPR’2023 T 43.21 54.10 26.90 25.10 7.30 23.50 21.70 3.60 1.90 1.60 4.10 24.40 8.10 24.20 1.60 1.10 0.00 13.10 6.60 5.70 13.41
H2GFormer-T [40] AAAI’2024 T 43.52 57.90 30.40 30.00 6.90 24.00 23.70 5.20 0.60 1.20 5.00 25.20 10.70 25.80 1.10 0.10 0.00 14.60 7.50 9.30 14.60
HASSC-T [39] CVPR’2024 T 42.87 55.30 29.60 25.90 11.30 23.10 23.00 2.90 1.90 1.50 4.90 24.80 9.80 26.50 1.40 3.00 0.00 14.30 7.00 7.10 14.38
SGN [25] TIP’2024 T 43.71 57.90 29.70 25.60 5.50 27.00 25.00 1.50 0.90 0.70 3.60 26.90 12.00 26.40 0.60 0.30 0.00 14.70 9.00 6.40 14.39
HTCL [15] ECCV’2024 T 44.23 64.40 34.80 33.80 12.40 25.90 27.30 5.70 1.80 2.20 5.40 25.30 10.80 31.20 1.10 3.10 0.90 21.10 9.00 8.30 17.09
Specifically, we follow the LSS view transformation paradigm and use depth bin assignment to project $F_{t}$, $F_{agg}$, and $M$into the 3D voxel space to obtain $V_{t}$, $V_{agg}$, and $V_{mask}$, respectively, as shown in Figure 5. First, we use $V_{mask}$ to distinguish occluded and non-occluded regions in the 3D voxel space. For the non-occluded region, we use the aggregate features $V_{agg}$ that fuse multiple cues. Subsequently, since the information from the corresponding position in the historical frame may be inaccurate due to occlusion, we use the voxel features from the current frame to update the occluded area to supplement the latest environmental information. Finally, by constructing a weighted matrix, we normalize the fused voxel features to ensure that there is no mutation at the boundary between the occluded and non-occluded areas, thereby improving the smoothness of the features. The specific operation is as follows:
@@
Finally, $V_{fine}$ enters the sparse voxel encoder for feature extraction, and then performs linear prediction to output dense semantic voxels $Y$.
@@ 3.6 Training Loss
In the FlowScene framework, we adopt the scene-class affinity loss $\mathcal{L}_{scal}$ from MonoScene [3] to optimize precision, recall, and specificity concurrently. The scene-class affinity loss is applied to semantic and geometric predictions, in conjunction with the cross-entropy loss weighted by class frequencies. Besides, the intermediate depth distribution for view transformation is supervised by the projections of LiDAR points, with the binary cross-entropy loss $\mathcal{L}_{d}$ following BEVDepth [18]. The overall loss function is formulated as follows:
@@
[…]
@@
Ours 70.01 58.81 46.98 29.83 4.44 3.78 16.71 8.71 7.77 60.70 16.99 39.59 6.01 43.17 9.45 37.32 25.14 17.35 18.12 10.63 7.56 19.12
Figure 6: Qualitative results on the SemanticKITTI validation set.
4 Experiments
To assess the effectiveness of our FlowScene, we conducted thorough experiments using the large outdoor datasets SemanticKITTI [1, 6], SSCBench-KITTI-360 [19, 21].
4.1 Experimental Setup
Datasets.
The SemanticKITTI[1, 6] dataset includes dense semantic scene completion annotations and labels a voxelized scene with 20 semantic classes. It consists of 10 training sequences, 1 validation sequence, and 11 testing sequences. RGB images are resized to $1280\times 384$ for input processing. The SSCBench-KITTI-360[19, 21] dataset contains 7 training sequences, 1 validation sequence, and 1 testing sequence, covering 19 semantic classes in total. The RGB images are resized to $1408\times 384$ for input processing.
Metrics.
Implementation Details.
We use RepVit [37] and FPN [22] to extract features for all images (The number of historical temporal frames n is set to 2). We use and freeze the GMFlow [43] optical flow estimation model to obtain optical flow information. We use the LSS paradigm for 2D-3D projection. The neighborhood cross-attention range is set to 7, and the number of attention heads is set to 8. Finally, the final outputs of SemantiKITTI is 20 classes, and SSCBench-KITTI-360 is 19 classes. All datasets have the scene size of $51.2m\times 51.2m\times 64m$ with the voxel grid size of $256\times 256\times 32$. By default, the model is trained for 25 epochs. We optimise the process, utilizing the AdamW optimizer with an initial learning rate of 1e-4 and a weight decay of 0.01. We also employ a multi-step scheduler to reduce the learning rate. All models are trained on two A100 Nvidia GPUs with 80G memory and batch size 4.
4.2 Main Results
Quantitative Results.
As shown in Table 1, we compare FlowScene with the latest public methods on the SemanticKITTI dataset, including approaches that use single-image input (S) and temporal image input (T). Temporal methods, such as VoxFormer [20], H2GFormer [40], HASSC [39], and SGN [25], utilize additional historical 5-frame input, while HTCL [15] uses a 3-frame historical input. In contrast, FlowScene uses only 2 historical frames as input, achieving the highest mIoU for the overall semantic metric and the highest IoU for the completion metric. Compared to the best-performing HTCL with temporal input, FlowScene improves the mIoU and IoU by 0.61% and 0.97%, respectively. When compared to the best CGFormer, which uses single-frame input, FlowScene achieves improvements of 1.07% in mIoU and 0.79% in IoU. Additionally, our method achieves the best or second-best results in most categories, outperforming or closely matching other methods. These results demonstrate the superiority of FlowScene in both geometry and semantics, effectively utilizing optical flow motion information and achieving temporal consistency.
@@ 4.2 Main Results
Due to the rich data samples and high-quality annotations in the SSCBench-KITTI-360 dataset, FlowScene outperforms current camera-based methods in both semantic and geometric analysis on the comprehensive SSCBench-KITTI-360 benchmark, as shown in Table 2. Our method achieves superior results in terms of both IoU and mIoU, surpassing all existing approaches.
Moreover, Table 3 illustrates the performance of FlowScene across three distance ranges (12.5m, 25.6m, 51.2m) on the SemanticKITTI validation set. It is evident that our approach significantly outperforms state-of-the-art methods at every tested distance.
Furthermore, as shown in Table 4, we compare the inference time and number of parameters of our method with other state-of-the-art methods on the SemanticKITTI validation set. FlowScene achieves state-of-the-art performance with a mIoU of 18.13%, while utilizing only 52.4M parameters. Additionally, FlowScene processes the extra 2-frame temporal image input with lower inference time, further demonstrating its efficiency and superior mIoU performance.
@@
Table 4: Comparison of inference time and number of parameters.
@@
Table 6: Ablation study of the different number of temporal inputs on the SemanticKITTI validation set.
Qualitative Visualizations.
4.3 Ablation Studies
We conduct extensive ablation experiments for FlowScene on the Semantickitti validation set. Specifically, we analyze the impact of different architecture component variations and temporal input in Table 6 and Table 5.
Optical Flow Estimation (OFE).
Flow-Guided Temporal Aggregation (FGTA).
Variants 3, 4, and 5 represent different configurations of the FGTA module: removing the FGTA module (Variant 3), removing the Occlusion Cross-Attention (Variant 4), and removing the Temporal Aggregation component (Variant 5). Variant 4 adaptively assigns weights to aggregate historical features, resulting in a 0.56% mIoU improvement (Variant 4 vs. Variant 3). Variant 5 uses Occlusion Cross-Attention to facilitate interaction between the current feature and the non-occluded areas in the historical frame, enhancing the texture and contextual information of the current frame’s features, further boosting the mIoU by 0.41% (Variant 5 vs. Variant 3).
Occlusion-Guided Voxel Refinement (OGVR).
Variant 6 represents the removal of the OGVR module, while Variant 7 uses convolution fusion to concatenate $V_{t}$ and $V_{agg}$. Even with this simple fusion strategy, a 0.45% mIoU improvement is achieved (Variant 7 vs. Variant 6). Variant 8 represents our final full model. Compared to Variant 7, the mask-based refinement strategy further improves the mIoU metric. It is worth noting that the OGVR module incurs no additional parameter overhead. Overall, compared to the baseline, our method achieves significant improvements in both completion and semantic metrics (+2.24% mIoU, +1.03% IoU).
Temporal Inputs.
As shown in Table 6, we evaluate the performance of temporal inputs with different numbers of frames. We observe that, as the number of frames increases, the time overhead also increases. However, the mIoU metric does not increase linearly, as the optical flow estimation model is less effective when the time interval between frames is long. As a result, inputs with 4 or 5 frames (t-4 and t-5) lead to reduced effectiveness. Considering both the experimental metrics and the time overhead, we use 2 frames as the input for our FlowScene method.

### v2

@@ 3.5 Occlusion-Guided Voxel Refinement
Specifically, as shown in Figure 5, we follow the LSS view transformation paradigm and use depth bin $D$ assignment to project $F_{t}$, $F_{agg}$, and $M$ into the 3D voxel space to obtain $V_{t}$, $V_{agg}$, and $V_{mask}$, respectively,
$V_{t}$ $=\operatorname{VoxelPooling}(F_{t}\otimes D),$
$V_{agg}$ $=\operatorname{VoxelPooling}(F_{agg}\otimes D),$ (6)
$V_{mask}$ $=\operatorname{VoxelPooling}(M\otimes D),$
where $\otimes$ is the dot product operation, VoxelPooling is the voxel pooling operation. First, we use $V_{mask}$ to distinguish occluded and non-occluded regions in the 3D voxel space. For the non-occluded region, we use the aggregate features $V_{agg}$ that fuse multiple cues. Subsequently, since the information from the corresponding position in the historical frame may be inaccurate due to occlusion, we use the voxel features from the current frame to update the occluded area to supplement the latest environmental information. Finally, by constructing a weighted matrix, we normalize the fused voxel features to ensure that there is no mutation at the boundary between the occluded and non-occluded areas, thereby improving the smoothness of the features. The specific operation is as follows:
@@
Finally, $V_{fine}$ enters the sparse voxel encoder for feature extraction, and then performs linear prediction to output dense semantic voxels $\mathbb{Y}$.
@@ 3.6 Training Loss
In the FlowScene framework, we adopt the scene-class affinity loss $\mathcal{L}_{scal}$ from MonoScene cao2022monoscene to optimize precision, recall, and specificity concurrently. The scene-class affinity loss is applied to semantic and geometric predictions, in conjunction with the cross-entropy loss weighted by class frequencies. Besides, the intermediate depth distribution for view transformation is supervised by the projections of LiDAR points, with the binary cross-entropy loss $\mathcal{L}_{d}$ following BEVDepth li2023bevdepth. The overall loss function is formulated as follows:
@@
4 Experiments
To assess the effectiveness of our FlowScene, we conducted thorough experiments using the large outdoor datasets SemanticKITTI behley2019semantickitti; Geiger2012kitti, SSCBench-KITTI-360 li2023sscbench; Liao2022kitti360. Information about datasets, metrics, and detailed implementation details is provided in the Technical Appendix, where additional experiments and analysis are also provided.
Methods Venues Input IoU
road
sidewalk
parking
other-grnd
building
car
truck
bicycle
motocycle
other-vehicle
vegetation
trunk
terrain
person
bicylist
motorcyclist
fence
pole
traf.-sign
mIoU
MonoScene cao2022monoscene CVPR’2022 S 34.16 54.70 27.10 24.80 5.70 14.40 18.80 3.30 0.50 0.70 4.40 14.90 2.40 19.50 1.00 1.40 0.40 11.10 3.30 2.10 11.08
TPVFormer huang2023tri CVPR’2023 S 34.25 55.10 27.20 27.40 6.50 14.80 19.20 3.70 1.00 0.50 2.30 13.90 2.60 20.40 1.10 2.40 0.30 11.00 2.90 1.50 11.26
OccFormer zhang2023occformer ICCV’2023 S 34.53 55.90 30.30 31.50 6.50 15.70 21.60 1.20 1.50 1.70 3.20 16.80 3.90 21.30 2.20 1.10 0.20 11.90 3.80 3.70 12.32
Symphonize jiang2024symphonize CVPR’2024 S 42.19 58.40 29.30 26.90 11.70 24.70 23.60 3.20 3.60 2.60 5.60 24.20 10.00 23.10 3.20 1.90 2.00 16.10 7.70 8.00 15.04
BRGScene li2023stereoscene IJCAI’2024 S 43.34 61.90 31.20 30.70 10.70 24.20 22.80 2.80 3.40 2.40 6.10 23.80 8.40 27.00 2.90 2.20 0.50 16.50 7.00 7.20 15.36
CGFormer CGFormer NIPS’2024 S 44.41 64.30 34.20 34.10 12.10 25.80 26.10 4.30 3.70 1.30 2.70 24.50 11.20 29.30 1.70 3.60 0.40 18.70 8.70 9.30 16.63
VoxFormer-T li2023voxformer CVPR’2023 T 43.21 54.10 26.90 25.10 7.30 23.50 21.70 3.60 1.90 1.60 4.10 24.40 8.10 24.20 1.60 1.10 0.00 13.10 6.60 5.70 13.41
H2GFormer-T wang2024h2gformer AAAI’2024 T 43.52 57.90 30.40 30.00 6.90 24.00 23.70 5.20 0.60 1.20 5.00 25.20 10.70 25.80 1.10 0.10 0.00 14.60 7.50 9.30 14.60
HASSC-T wang2024HASSC CVPR’2024 T 42.87 55.30 29.60 25.90 11.30 23.10 23.00 2.90 1.90 1.50 4.90 24.80 9.80 26.50 1.40 3.00 0.00 14.30 7.00 7.10 14.38
SGN mei2024sgn TIP’2024 T 43.71 57.90 29.70 25.60 5.50 27.00 25.00 1.50 0.90 0.70 3.60 26.90 12.00 26.40 0.60 0.30 0.00 14.70 9.00 6.40 14.39
HTCL li2024htcl ECCV’2024 T 44.23 64.40 34.80 33.80 12.40 25.90 27.30 5.70 1.80 2.20 5.40 25.30 10.80 31.20 1.10 3.10 0.90 21.10 9.00 8.30 17.09
4.1 Main Results
Quantitative Results.
As shown in Table 1, we compare FlowScene with the latest public methods on the SemanticKITTI dataset, including approaches that use single-image input (S) and temporal image input (T). Temporal methods, such as VoxFormer li2023voxformer, H2GFormer wang2024h2gformer, HASSC wang2024HASSC, and SGN mei2024sgn, utilize additional historical 5-frame input, while HTCL li2024htcl uses a 3-frame historical input. In contrast, FlowScene uses only 2 historical frames as input, achieving the highest mIoU for the overall semantic metric and the highest IoU for the completion metric. Compared to the best-performing HTCL with temporal input, FlowScene improves the mIoU and IoU by 0.61% and 0.97%, respectively. When compared to the best CGFormer, which uses single-frame input, FlowScene achieves improvements of 1.07% in mIoU and 0.79% in IoU. Additionally, our method achieves the best or second-best results in most categories, outperforming or closely matching other methods. These results demonstrate the superiority of FlowScene in both geometry and semantics, effectively utilizing optical flow motion information and achieving temporal consistency.
@@ 4.1 Main Results
car
bicycle
motocycle
truck
other-vehicle
person
road
parking
sidewalk
other-grnd
building
fence
vegetation
terrain
pole
traf.-sign
other-struct.
other-obj.
@@
CGFormer - - 48.07 29.85 3.42 3.96 17.59 6.79 6.63 63.85 17.15 40.72 5.53 42.73 8.22 38.80 24.94 16.24 17.45 10.18 6.77 20.05
Ours 69.70 58.99 46.95 32.48 1.87 4.93 25.47 14.86 9.62 60.53 16.49 36.13 8.58 39.66 9.62 39.82 13.32 17.52 14.35 16.25 13.08 20.81
To demonstrate the diversity of our model, we conducted experiments on the SSCBench-KITTI-360 dataset, as shown in Table 2. It is worth noting that our method has a huge advantage in the performance of potential moving objects (such as car, truck, other-vehicle, person, etc.)
Moreover, Table 3 illustrates the performance of FlowScene across three distance ranges (12.5m, 25.6m, 51.2m) on the SemanticKITTI validation set. It is evident that our approach significantly outperforms state-of-the-art methods at every tested distance. Furthermore, as shown in Table 4, we compare the inference time and number of parameters of our method with other state-of-the-art methods on the SemanticKITTI validation set. The inference time of MonoScene is optimal because of its FLoSP feature projection method. But, FlowScene achieves state-of-the-art performance with a mIoU of 18.13%, while utilizing only 52.4M parameters. Additionally, FlowScene processes the extra 2-frame temporal image input with lower inference time, further demonstrating its efficiency and superior mIoU performance.
@@
Table 4: Comparison of inference time and number of parameters.
@@
Qualitative Visualizations.
Figure 6: Qualitative results on the SemanticKITTI validation set.
Table 5: Analysis of Static and Dynamic Objects
Method SemanticKITTI SSCBench-KITTI-360
Dynamic Static All-mIoU Dynamic Static All-mIoU
MonoScene 3.81 16.36 11.08 5.21 15.87 12.13
VoxFormer 4.45 19.91 13.41 4.69 12.19 11.91
Symphonie 5.71 21.83 15.04 13.85 20.94 18.58
CGFormer 5.48 24.75 16.63 11.37 24.38 20.05
Ours 7.50 25.29 17.70 14.87 23.78 20.81
Figure 7: Dynamic object modeling case visualization results.
4.2 Analysis of Static and Dynamic Objects
To better understand the performance of our method across different types of semantic categories, we conduct a detailed analysis by separating static and dynamic object classes in the autonomous driving datasets. For SemanticKITTI, we define dynamic objects as the classes that are likely to involve motion:   car,   truck,   bicycle,   motocycle,   other-vehicle,   person and   bicylist as dynamic objects and others as static objects. For SSCBench-KITTI-360, we classified   car,   bicycle,   motocycle,   truck,   other-vehicle and   person as dynamic objects and others as static objects. Table 12 presents a comparative evaluation of our method against several prior state-of-the-art approaches, reporting mean IoU (mIoU) separately for dynamic objects, static objects, and the overall average.
Our method clearly outperforms all baselines in both datasets. Notably: On SemanticKITTI, our model achieves 7.50% mIoU for dynamic objects—the highest among all methods, surpassing CGFormer (5.48%) and Symphonie (5.71%). For static objects, we also lead with 25.29%, again outperforming CGFormer (24.75%). On SSCBench-KITTI-360, our method obtains 14.87% mIoU on dynamic objects, significantly ahead of CGFormer (11.37%) and Symphonie (13.85%). Although CGFormer slightly surpasses us on static objects (24.38% vs. 23.78%), our method achieves the highest overall score of 20.81% mIoU. These results demonstrate that FlowScene plays a crucial role in handling motion and temporal variation, making it especially effective in recognizing and completing dynamic objects. Unlike previous approaches that rely on frame stacking or static assumptions, our model adapts to motion patterns, improving semantic consistency across time.
Figure 7 illustrates the qualitative results of dynamic object modeling in a challenging real-world scenario. As shown, the cyclist in motion is effectively captured by the optical flow module, which accurately estimates the movement across frames. By leveraging this motion information, our model performs kinematic compensation, aligning temporal features and preserving object structure throughout the sequence. This results in a more coherent and complete 3D semantic reconstruction of the dynamic scene. To further support this case, we provide a supplementary video that offers an intuitive, frame-by-frame visualization of the temporal alignment and dynamic object modeling. This additional material highlights the superior capability of our method to handle complex motion compared to static-assumption baselines.
@@ 4.2 Analysis of Static and Dynamic Objects
4.3 Ablation Studies
We conduct extensive ablation experiments for FlowScene on the Semantickitti validation set. Specifically, we analyze the impact of different architecture component variations in Table 6.
Optical Flow Estimation (OFE).
Flow-Guided Temporal Aggregation (FGTA).
In Table 6, Variants 3, 4, and 5 represent different configurations of the FGTA module: removing the FGTA module (Variant 3), removing the Occlusion Cross-Attention (Variant 4), and removing the Temporal Aggregation component (Variant 5). Variant 4 adaptively assigns weights to aggregate historical features, resulting in a 0.56% mIoU improvement (Variant 4 vs. Variant 3). Variant 5 uses Occlusion Cross-Attention to facilitate interaction between the current feature and the non-occluded areas in the historical frame, enhancing the texture and contextual information of the current frame’s features, further boosting the mIoU by 0.41% (Variant 5 vs. Variant 3).
Occlusion-Guided Voxel Refinement (OGVR).
In Table 6, Variant 6 represents the removal of the OGVR module, while Variant 7 uses convolution fusion to concatenate $V_{t}$ and $V_{agg}$. Even with this simple fusion strategy, a 0.45% mIoU improvement is achieved (Variant 7 vs. Variant 6). Variant 8 represents our final full model. Compared to Variant 7, the mask-based refinement strategy further improves the mIoU metric. It is worth noting that the OGVR module incurs no additional parameter overhead.
Table 7: Ablation study for temporal alignment strategy.
Method IoU(%) mIoU(%)
Stack 43.98 15.89
VoxFormer-T li2023voxformer 44.15 13.35
HTCL li2024htcl 45.51 17.13
Flow-Guided [Ours] 45.01 18.13
Table 8: Ablation study for different number of temporal inputs.
@@ 4.3 Ablation Studies
Table 9: Ablation study for optical flow networks.
Method IoU(%) mIoU(%) Params(M)
PWC-Net+ sun2018pwc 43.31 17.13 8.8
RAFT teed2020raft 44.12 17.56 5.3
FlowFormer huang2022flowformer 44.33 17.74 18.2
GMFlow xu2022gmflow 45.01 18.13 4.7
Table 10: Ablation study for image backbone networks. Acc represents the classification accuracy of each pre-trained model on ImageNet russakovsky2015imagenet.
Method IoU(%) mIoU(%) Params(M) Acc(%)
ResNet50 he2016resnet 44.12 16.98 25.6 79.3
EfficientNetB7 tan2019efficientnet 44.31 17.63 63.8 84.4
RepVit-M2.3 wang2023repvit 45.01 18.13 22.4 83.3
Overall, compared to the baseline, our method achieves significant improvements in both completion and semantic metrics (+2.24% mIoU, +1.03% IoU).
Temporal Alignment Strategy and Temporal Inputs.
Table 7 presents the results of an ablation study comparing different temporal alignment strategies used for fusing features across frames. Our Flow-Guided strategy delivers the best mIoU of 18.13%, showing that optical flow-guided alignment is highly effective for preserving fine-grained semantic consistency, particularly for dynamic scenes. As shown in Table 8, we evaluate the performance of temporal inputs with different numbers of frames. We observe that, as the number of frames increases, the time overhead also increases. However, the mIoU metric does not grow linearly, as the quality of optical flow prediction decreases when the time interval between frames is longer. As a result, inputs with 4 or 5 frames (t-4 and t-5) lead to reduced effectiveness. Considering both the experimental metrics and the time overhead, we use 2 frames as the input for our FlowScene method.
Optical Flow Networks.
Table 9 presents the performance of different optical flow networks. We compare several state-of-the-art methods, including PWC-Net sun2018pwc, RAFT teed2020raft, and FlowFormer huang2022flowformer, along with our setting, GMFlow xu2022gmflow, which is highlighted in the last row. Our setting achieves the highest IoU of 45.01% and mIoU of 18.13%, outperforming all other methods in both metrics. These results suggest that GMFlow effectively captures motion cues and integrates them into the semantic scene completion task, providing superior performance over the other optical flow networks tested, with significantly fewer parameters.
Image Backbone Networks.
Table 10 examines the impact of different backbone networks on the performance of FlowScene. The study compares EfficientNetB7 tan2019efficientnet, ResNet50 he2016resnet, and RepVit-M2.3 wang2023repvit(our setting). Our method, using RepVit-M2.3, achieves the highest IoU of 45.01% and mIoU of 18.13%, surpassing both EfficientNetB7 (44.31% IoU, 17.63% mIoU) and ResNet50 (44.12% IoU, 16.98% mIoU). RepVit-M2.3, though achieving the best performance, maintains a relatively low parameter count of 22.4M. In comparison, EfficientNetB7 has a much higher parameter count of 63.8M, while ResNet50 is more parameter-efficient at 25.6M. RepVit-M2.3 offers a good balance between performance and parameter count, making it an ideal choice for our backbone network. Different image encoders have significant improvements over the baseline, demonstrating the effectiveness of our entire model.